# Self-Supervised Terrain Representation Learning from Unconstrained Robot Experience

### v1

@@ Abstract
Terrain awareness, i.e., the ability to identify and distinguish different types of terrain, is a critical ability that robots must have to succeed at autonomous off-road navigation. Current approaches that provide robots with this awareness either rely on labeled data which is expensive to collect, engineered features and cost functions that may not generalize, or expert human demonstrations which may not be available. Towards endowing robots with terrain awareness without these limitations, we introduce Self-supervised TErrain Representation LearnING (sterling), a novel approach for learning terrain representations that relies solely on easy-to-collect, unconstrained (e.g., non-expert), and unlabelled robot experience, with no additional constraints on data collection. sterling employs a novel multi-modal self-supervision objective through non-contrastive representation learning to learn relevant terrain representations for terrain-aware navigation. Through physical robot experiments in off-road environments, we evaluate sterling features on the task of preference-aligned visual navigation and find that sterling features perform on par with fully-supervised approaches and outperform other state-of-the-art methods with respect to preference alignment. Additionally, we perform a large-scale experiment of semi-autonomously hiking a 3-mile long trail which sterling completes successfully with only two manual interventions, demonstrating robustness to real-world off-road conditions.
11footnotetext: The University of Texas at Austin, Austin, TX, USA22footnotetext: Cornell University, Ithaca, NY, USA33footnotetext: Army Research Laboratory, Austin, TX, USA44footnotetext: Sony AI, North America, Austin, TX, USA
@@ 1 Introduction
Off-road navigation is emerging as a crucial capability for autonomous mobile robots envisioned for use in a growing number of outdoor applications such as agricultural operations [1], package delivery [2], and search and rescue [3]. The ability to visually identify distinct terrain features that are relevant to the navigation task, referred to here as terrain awareness, is necessary for enabling off-road navigation.
Endowing robots with terrain awareness is a challenging problem in autonomous off-road navigation [4, 5, 6]. Prior work on terrain awareness has typically relied on difficult-to-collect curated datasets [7, 8, 9, 10] or has been focused on particular tasks [11, 12, 13, 14, 15, 4] and is not amenable to downstream task changes [16, 15, 17]. To enable off-road autonomy at scale, new approaches are needed that address the perception problem of learning relevant terrain representations in a self-supervised way from easy-to-collect unconstrained robot experiences gathered without requiring an expert agent in the field. Moreover, the learned representations need to be general enough to support a variety of downstream navigation tasks (e.g., changing preferences over terrain types). Learning such rich visual features of the terrain through self-supervision by leveraging multi-modal sensor data collected onboard a mobile robot can help tackle many off-road navigation problems such as high-speed navigation [4, 18], traversing challenging terrain [19, 20, 21, 22, 23, 24], and preference-aligned planning [25, 10, 26, 27, 28].
@@
Towards robot terrain awareness from unconstrained and unlabelled robot experience, we introduce Self-supervised TErrain Representation LearnING (sterling) [^1] A preliminary version of this work was presented at the PT4R workshop at ICRA 2023 [29], a novel approach to learning terrain representations for off-road navigation. sterling utilizes easy-to-collect unconstrained robot experiences, without any additional restrictions on data collection (e.g., that the trajectories be of expert demonstrations), and employs a novel multi-modal self-supervision objective through non-contrastive, non-reconstructive representation learning to learn relevant terrain representations useful for visual terrain-aware navigation. The key contributions of this paper are— 1) Self-supervised TErrain Representation LearnING (sterling), a novel approach that learns terrain representations from easy-to-collect unconstrained robot experiences, 2) Detailed evaluation of sterling against baseline methods on the task of operator preference-aligned off-road navigation, and 3) A large-scale qualitative experiment of semi-autonomously hiking a 3-mile long trail, demonstrating the effectiveness of sterling-features.
We execute all of our off-road physical robot experiments on a Boston Dynamics Spot and compare features learned using sterling to existing visual representation learning-based methods se-r [5] and ganav [10], self-supervised cost learning approach rca [17], and a fully-supervised oracle. We find that sterling performs on par with or better than baseline approaches in all the outdoor environments with respect to preference alignment, without requiring expert labels or demonstrations.
@@ 2.1 Supervised Methods
End-to-End Learning: The initial success of applying learning-based solutions to off-road terrain-aware navigation was by LeCun et al. [30] who used a convolutional network to learn to drive in off-road conditions. More recently, Bojarski et al. [11] trained a deep neural network end-to-end using several miles of driving data collected on a vehicle in the real world. While both approaches were promising in urban and off-road environments, end-to-end methods require large amounts of data and are well-known to suffer from domain and covariate shifts [31, 32].
Image Segmentation: Unlike end-to-end approaches that learn behaviors, segmentation-based approaches either classify or pixel-wise segment terrains into known classes [33, 10, 34], trained using supervised learning on large labeled datasets. Guan et al. [10] propose ganav, a transformer-based architecture to pixel-wise segment terrains, trained on RELLIS [8] and RUGD [7] datasets, with manually assigned terrain costs. While effective at terrain awareness, segmentation-based methods are fixed to the specific terrain types available in the datasets and require additional labeling effort to generalize to novel terrains. In sterling, we do not require semantically labeled datasets and learn terrain representations from unconstrained experience collected onboard a mobile robot.
@@ 2.2 Self-Supervised Learning
Cost Learning: Several methods have applied self-supervision to assign traversability costs for the downstream off-road navigation task [17, 39, 16, 40]. Specifically, these methods rely on inertial spectral features [17], future predictive models [16], inertial-odometry errors [39], or force-torque values from foothold positions [40, 41] as self-supervision signals to learn a traversability cost map, used to evaluate candidate actions. More recently, Frey et al. [42] have proposed an online traversability estimation approach inspired by the above self-supervision schemes. Instead of inferring costs or rewards using self-supervision for a fixed task, in this work, we focus on learning relevant visual features from unconstrained robot experiences that could be used in downstream tasks. This framework allows a designer to reuse features across tasks without retraining entirely from scratch.
@@ 3 Approach
In this section, we introduce the self-supervised terrain representation learning approach, sterling, proposed in this work. We first describe the offline pre-processing performed on recorded unconstrained robot data and then summarize the self-supervision objectives. Finally, we describe the problem formulation for preference-aligned off-road navigation and present how features learned using sterling can be utilized within a planner for terrain-aware and preference-aligned navigation.
@@ 3.1 Data-Collection and Pre-Processing
sterling learns terrain representations from unconstrained, unlabeled robot experiences collected using any navigation policy. This policy may be, for instance, non-expert human teleoperation, curiosity-driven exploration, or point-to-point navigation using any underlying planner. Compared to requiring a human expert to provide teleoperated demonstrations and labels, collecting this type of robot experience is cheap and easy. We additionally assume that the robot is equipped with multiple sensors, e.g., an egocentric RGB camera, odometry sensors, onboard IMU, proprioceptive, and tactile sensors, that together provide rich multi-modal observations as the robot traverses over different terrains collecting experience. sterling leverages this multi-modal data by using the correlation between different modalities to inform the learned terrain representations.
@@
Visual Patch Extraction: The egocentric camera observations are homography-projected into a virtual bird’s eye view (BEV) frame, assuming that the ground is a flat plane, using the intrinsic and extrinsic camera matrices. As shown in Fig. 1, we project the robot’s trajectory onto the bird’s eye view frame and extract 64-by-64 pixels (equivalent to the robot’s footprint of 0.5-by-0.5 meters) square visual patches of the terrain along with the corresponding inertial, proprioceptive, and tactile measurements at the same location, along the trajectory. Since the terrain at a particular location $s_{k}$ is unobservable when the robot itself is at $s_{k}$ (i.e., it is underneath the robot), we extract terrain image patches corresponding to $s_{k}$ from BEV observations made by the robot from previous locations $s_{k-1},s_{k-2},\ldots$ along its trajectory. Fig. 1 illustrates the offline patch extraction process from two previous viewpoints, however, we extract patches from up to 20 previous viewpoints within 2 meters. Although just one viewpoint is sufficient to learn the correlation between visual and other sensor observations, when planning to navigate, the robot will need to visually evaluate terrain at future locations, and therefore sterling also seeks representations that are invariant to patch differences due to viewpoint, also known as viewpoint invariance.
@@ 3.2 Non-Contrastive Terrain Representation Learning
Fig. 2 provides an overview of the self-supervised representation learning framework adopted in sterling. A parameterized visual encoder (4-layer CNN with 0.25 million parameters) encodes terrain image patch observations $v_{1}$ and $v_{2}$ of the same location $s$ into visual representations $\phi_{v_{1}}$ and $\phi_{v_{2}}$, respectively. Similarly, an inertial-proprio-tactile encoder (4-layer MLP with 0.25 million parameters) encodes frequency domain ipt observations of the robot at that location to an inertial-proprio-tactile representation $\phi_{i}$. We follow the framework of prior self-supervised representation learning algorithms from the computer vision community such as vicreg [38], and utilize a parameterized projector network (2-layer MLP with 0.25 million parameters) that maps encoded visual and non-visual representations independently to a higher-dimensional feature space $\psi_{v_{1,2}}$ and $\psi_{i}$ respectively, over which the self-supervision objectives are computed. The sterling objective composed of the multi-modal correlation $\mathcal{L}_{MM}(\psi_{v_{1,2}},\psi_{i})$ and viewpoint-invariance $\mathcal{L}_{VI}(\psi_{v_{1}},\psi_{v_{2}})$ objectives are defined as:
@@
$\mathcal{L}_{\textsc{vicr}eg{}}$ is the vicreg loss that is composed of variance-invariance-covariance representation learning objectives, as proposed by Bardes et al. [38]. Given two alternate projected representations $Z$ and $Z^{\prime}$ of a data sample (in sterling, $Z$ and $Z^{\prime}$ are projected representations of the visual and non-visual sensor modalities), the vicreg loss is defined as $\mathcal{L}_{\textsc{vicr}eg{}}(Z,Z^{\prime})=\lambda s(Z,Z^{\prime})+\mu[v(Z)+v(Z^{\prime})]+\nu[c(Z)+c(Z^{\prime})]$. $\lambda$, $\mu$, and $\nu$ are hyper-parameters and the functions $v$, $s$, and $c$ are the variance, invariance, and covariance terms computed on a mini-batch of projected features. More details on the individual terms in the loss function is provided in Sec. 9.5. On a mini-batch of data containing paired terrain image patches and ipt observations, we compute the $\mathcal{L}_{\textsc{sterling}{}}$ loss and update parameters of the two encoder networks and the shared projector network together using Adam optimizer.
@@
In all experiments, we use a legged Boston Dynamics Spot robot and collect robot experiences on eight different types of terrain around the university campus that we labeled as mulch, pebble sidewalk, cement sidewalk, grass, bushes, marbled rock, yellow bricks, and red bricks. We refer the reader to Supplemental Sec. 7 for more details on data collection for this work.
@@ 3.3 Preference-Aligned Off-Road Navigation
In this subsection, we describe the problem formulation of terrain-aware navigation and how sterling features can be utilized within a planner to navigate in a preference-aligned manner. While the main novelty of our work is on learning terrain representations from unconstrained robot experience, we use preference-aligned terrain-aware navigation as a real-world task to evaluate sterling features.
@@ 4.1 Evaluating Terrain-Awareness via Robot Experiments
Fig. 3 shows the operator’s (first author) terrain preferences for all Envs. 1 to 5, and the performance of baseline approaches, including an operator-demonstrated trajectory for reference. In all environments, we see that sterling navigates in a terrain-aware manner while adhering to the operator’s preferences. Note that although Fully-Supervised also completes the task successfully, it requires privileged information such as terrain labels during training, whereas sterling does not require such supervision, and can potentially be used on large datasets containing unlabelled, unconstrained robot experiences. ganav, trained on the rugd dataset fails to generalize to unseen real-world conditions. rca uses inertial spectral features to learn terrain traversability costs and hence does not adhere to operator preference. se-r does not address viewpoint invariance which is a significant problem in vision-based off-road navigation and hence performs poorly in Envs. 1 and 2. We perform additional experiments in an outdoor environment (Env. 6) to study adherence to operator preferences, detailed in Supplementary Sec. 9.1.
@@ 6 Conclusion
In this paper, we introduce Self-supervised TErrain Representation LearnING (sterling), a novel framework for learning terrain representations from easy-to-collect, unconstrained (e.g., non-expert), and unlabelled robot experience. sterling utilizes non-contrastive representation learning through viewpoint invariance and multi-modal correlation self-supervision objectives to learn relevant terrain representations for visual navigation. We show how features learned through sterling can be utilized to learn operator preferences over terrains and integrated within a planner for preference-aligned navigation. We evaluate sterling against state-of-the-art alternatives on the task of preference-aligned visual navigation on a Spot robot and find that sterling outperforms other methods and performs on par with a fully-supervised baseline. We additionally perform a qualitative large-scale experiment by successfully hiking a 3-mile-long trail using sterling, demonstrating its robustness to off-road conditions in the real world.

### v2

# STERLING: Self-Supervised Terrain Representation Learning from Unconstrained Robot Experience

@@ Abstract
Terrain awareness, i.e., the ability to identify and distinguish different types of terrain, is a critical ability that robots must have to succeed at autonomous off-road navigation. Current approaches that provide robots with this awareness either rely on labeled data which is expensive to collect, engineered features and cost functions that may not generalize, or expert human demonstrations which may not be available. Towards endowing robots with terrain awareness without these limitations, we introduce Self-supervised TErrain Representation LearnING (sterling), a novel approach for learning terrain representations that relies solely on easy-to-collect, unconstrained (e.g., non-expert), and unlabeled robot experience, with no additional constraints on data collection. sterling employs a novel multi-modal self-supervision objective through non-contrastive representation learning to learn relevant terrain representations for terrain-aware navigation. Through physical robot experiments in off-road environments, we evaluate sterling features on the task of preference-aligned visual navigation and find that sterling features perform on par with fully-supervised approaches and outperform other state-of-the-art methods with respect to preference alignment. Additionally, we perform a large-scale experiment of semi-autonomously hiking a 3-mile long trail which sterling completes successfully with only two manual interventions, demonstrating robustness to real-world off-road conditions. Robot experiment videos and more details can be found in the appendix and the project website https://hareshkarnan.github.io/sterling/
11footnotetext: The University of Texas at Austin, Austin, TX, USA22footnotetext: Cornell University, Ithaca, NY, USA33footnotetext: Army Research Laboratory, Austin, TX, USA44footnotetext: Sony AI, North America
@@ 1 Introduction
Off-road navigation is emerging as a crucial capability for autonomous mobile robots envisioned for use in a growing number of outdoor applications such as agricultural operations [1], package delivery [2], and search and rescue [3]. Endowing robots with this capability has, however, proved to be challenging and remains an active area of research.
One particularly difficult challenge in off-road autonomous navigation is that of providing the robot with terrain awareness, i.e., the ability to identify distinct terrain features that are relevant to a wide variety of downstream tasks (e.g., changing preferences over terrain types) [4, 5, 6, 7, 8, 9, 10, 11]. While a litany of prior work has attempted to address this challenge [12, 13, 14, 15], existing approaches typically rely on difficult-to-collect curated datasets [16, 17, 18, 19, 20] or has been focused on particular tasks [21, 22, 23, 24, 25, 9] and is not amenable to downstream task changes [26, 25, 7]. These limitations prevent existing approaches from appropriately scaling to the vast distribution of terrains and navigation tasks in the real world.
Toward overcoming the scalability challenges in terrain awareness, we introduce Self-supervised TErrain Representation LearnING (sterling) [^1] A preliminary version of this work was presented at the PT4R workshop at ICRA 2023 [27], a novel approach to learning terrain representations for off-road navigation. sterling learns an encoding function that maps high-dimensional, multi-modal sensor data to low-dimensional, terrain-aware representations that amplify differences important for navigation and attenuate differences due to extraneous factors such as changes in viewpoint and lighting. Importantly, sterling works with easy-to-collect unconstrained and unlabeled robot data, thereby providing a scalable pathway to data collection and system improvement for the wide variety of terrain and downstream tasks that off-road robots must face.
To evaluate sterling, we apply it to the problem of preference-aligned off-road navigation and provide a detailed comparison to existing approaches to this problem, including rca [7], ganav [19], se-r [8], and a fully-supervised oracle. We find that sterling enables performance on par with or better than these existing approaches without requiring any expert labels or demonstrations. Additionally, we report the results of a large-scale qualitative experiment in which sterling enabled semi-autonomous robot navigation on a 3-mile long hiking trail.
@@
The key contributions of this paper are— 1) Self-supervised TErrain Representation LearnING (sterling), a novel approach that learns terrain representations from easy-to-collect unconstrained robot experiences, 2) Detailed evaluation of sterling against baseline methods on the task of operator preference-aligned off-road navigation, and 3) A large-scale qualitative experiment of semi-autonomously hiking a 3-mile long trail, demonstrating the effectiveness of sterling-features.
@@ 2.1 Supervised Methods
End-to-End Learning: The initial success of applying learning-based solutions to off-road terrain-aware navigation was by LeCun et al. [28] who used a convolutional network to learn to drive in off-road conditions. More recently, Bojarski et al. [21] trained a deep neural network end-to-end using several miles of driving data collected on a vehicle in the real world. While both approaches were promising in urban and off-road environments, end-to-end methods require large amounts of data and are well-known to suffer from domain and covariate shifts [29, 30, 31].
Image Segmentation: Unlike end-to-end approaches that learn behaviors, segmentation-based approaches seek to characterize terrain using a set of known semantic classes, and the resulting semantic features are consumed by downstream planning and control techniques for navigation [32, 19, 33]. Guan et al. [19] propose ganav, a transformer-based architecture to pixel-wise segment terrains, trained on RELLIS [17] and RUGD [16] datasets, with manually assigned terrain costs. While effective at terrain awareness, segmentation-based methods are fixed to the specific terrain types available in the datasets and require additional labeling effort to generalize to novel terrains. In sterling, we do not require semantically labeled datasets and learn terrain representations from unconstrained experience collected onboard a mobile robot.
@@ 2.2 Self-Supervised Learning
Cost Learning: Several methods have applied self-supervision to assign traversability costs for the downstream off-road navigation task [7, 38, 26, 39, 40, 41, 42]. Specifically, these methods rely on inertial spectral features [7], future predictive models [26], inertial-odometry errors [38], or force-torque values from foothold positions [39, 43] as self-supervision signals to learn a traversability cost map, used to evaluate candidate actions. More recently, Frey et al. [44] have proposed an online traversability estimation approach inspired by the above self-supervision schemes. Instead of inferring costs or rewards using self-supervision for a fixed task, in this work, we focus on learning relevant visual features from unconstrained robot experiences that could be used in downstream tasks. This framework allows a designer to reuse features across tasks without retraining entirely from scratch.
@@ 3 Approach
In this section, we introduce the self-supervised terrain representation learning approach, sterling, proposed in this work. We first describe the offline pre-processing performed on unconstrained robot data and then summarize the self-supervision objectives. Finally, we describe the problem formulation for preference-aligned off-road navigation and present how features learned using sterling can be utilized within a planner for terrain-aware and preference-aligned navigation.
@@ 3.1 Data-Collection and Pre-Processing
sterling learns terrain representations from unconstrained, unlabeled robot experiences collected using any navigation policy. This policy may be, for instance, non-expert human teleoperation, curiosity-driven exploration [45], or point-to-point navigation using any underlying planner. Compared to requiring a human expert to provide teleoperated demonstrations and labels, collecting this type of robot experience is cheap and easy, thereby providing a scalable pathway to data collection and system improvement. We additionally assume that the robot is equipped with multiple sensors, e.g., an egocentric RGB camera, odometry sensors, onboard IMU, proprioceptive, and tactile sensors, that together provide rich multi-modal observations as the robot traverses over different terrains collecting experience. sterling leverages this multi-modal data by using the correlation between different modalities to inform the learned terrain representations.
@@
Visual Patch Extraction: The egocentric camera observations are homography-projected into a virtual bird’s eye view (BEV) frame, assuming that the ground is a flat plane, using the intrinsic and extrinsic camera matrices. As shown in Fig. 1, we project the robot’s trajectory onto the BEV frame and extract 64-by-64 pixels (equivalent to the robot’s footprint of 0.5-by-0.5 meters) square visual patches of terrain along with the corresponding inertial, proprioceptive, and tactile observations at the same location, along the trajectory. Since the terrain at $s_{k}$ is unobservable when the robot itself is at $s_{k}$ (i.e., it is underneath the robot), we extract terrain image patches corresponding to $s_{k}$ from BEV observations at previous locations $s_{k-1},s_{k-2},\ldots$ along its trajectory. Fig. 1 illustrates the offline patch extraction process from two previous viewpoints, however, we extract patches from up to 20 previous viewpoints within 2 meters. Although just one viewpoint is sufficient to learn the correlation between visual and other sensor observations, when planning to navigate, the robot will need to visually evaluate terrain at future locations, and therefore sterling also seeks representations that are invariant to patch differences due to viewpoint, also known as viewpoint invariance.
@@ 3.2 Non-Contrastive Terrain Representation Learning
Fig. 2 provides an overview of the self-supervised representation learning framework adopted in sterling. A parameterized visual encoder (4-layer CNN with 0.25 million parameters) encodes terrain image patch observations $v_{1}$ and $v_{2}$ of the same location $s$ into visual representations $\phi_{v_{1}}$ and $\phi_{v_{2}}$, respectively, collectively referred to as $\phi_{v_{1,2}}$ for brevity. Similarly, an inertial-proprio-tactile encoder (4-layer MLP with 0.25 million parameters) encodes frequency domain ipt observations of the robot at that location to an inertial-proprio-tactile representation $\phi_{i}$. We follow the framework of prior self-supervised representation learning algorithms from the computer vision community such as vicreg [37], and utilize a parameterized projector network (2-layer MLP with 0.25 million parameters) that maps encoded visual and non-visual representations independently to a higher-dimensional feature space $\psi_{v_{1,2}}$ and $\psi_{i}$ respectively, over which the self-supervision objectives are computed. The sterling objective composed of the multi-modal correlation $\mathcal{L}_{MM}(\psi_{v_{1,2}},\psi_{i})$ and viewpoint-invariance $\mathcal{L}_{VI}(\psi_{v_{1}},\psi_{v_{2}})$ objectives are defined as:
@@
$\mathcal{L}_{\textsc{vicr}eg{}}$ is the vicreg loss that is composed of variance-invariance-covariance representation learning objectives, as proposed by Bardes et al. [37]. Given two alternate projected representations $Z$ and $Z^{\prime}$ of a data sample (in sterling, $Z$ and $Z^{\prime}$ are projected representations of the visual and non-visual sensor modalities), the vicreg loss is defined as $\mathcal{L}_{\textsc{vicr}eg{}}(Z,Z^{\prime})=\lambda s(Z,Z^{\prime})+\mu[v(Z)+v(Z^{\prime})]+\nu[c(Z)+c(Z^{\prime})]$. Note that while Bardes et al. use vicreg to learn representations from visual inputs using artificial image augmentations, in this work, we extend vicreg to multi-modal inputs and use real-world augmentations via multi-viewpoint image patches as described in Sec. 3.1. $\lambda$, $\mu$, and $\nu$ are hyper-parameters and the functions $v$, $s$, and $c$ are the variance, invariance, and covariance terms computed on a mini-batch of projected features. We refer the reader to Bardes et al. [37] for additional details on the individual terms and also define them here for completeness. The variance term $v$ is a hinge function defined as $v(Z)=\frac{1}{d}\sum\limits_{j=1}^{d}max(0,\gamma-S(z^{j},\epsilon))$, where $S$ is the standard deviation, and $d$ is the dimensionality of the projected feature space. $c$ is the covariance term, defined as $c(Z)=\frac{1}{d}\sum\limits_{i\neq j}[C(Z)]^{2}_{i,j}$, where $C(Z)$ is the covariance matrix of $Z$. $s$ is the invariance term defined as $s(Z,Z^{{}^{\prime}})=\frac{1}{n}\sum\limits_{i}||z_{i}-z^{{}^{\prime}}_{i}||$. More details on the individual terms in the loss function are provided in Sec. 9.5. We apply an $l^{2}$ norm on the visual and non-visual features to ensure they are on a hypersphere, which helped improve the quality of learned representations. On a mini-batch of data containing paired terrain image patches and ipt observations, we compute the $\mathcal{L}_{\textsc{sterling}{}}$ loss and update parameters of the two encoder networks and the shared projector network together using Adam optimizer.
@@ 3.3 Preference-Aligned Off-Road Navigation
In this subsection, we describe the downstream navigation task of preference-aligned visual navigation that we focus on when evaluating sterling.
@@ 4.1 Evaluating Terrain-Awareness via Robot Experiments
Fig. 3 shows the operator’s (first author) terrain preferences for all Envs. 1 to 5, and the performance of baseline approaches, including an operator-demonstrated trajectory for reference. In all environments, we see that sterling navigates in a terrain-aware manner while adhering to the operator’s preferences. Note that although Fully-Supervised also completes the task successfully, it requires privileged information such as terrain labels during training, whereas sterling does not require such supervision, and can potentially be used on large datasets containing unlabeled, unconstrained robot experiences. ganav, trained on the rugd dataset fails to generalize to unseen real-world conditions. rca uses inertial spectral features to learn terrain traversability costs and hence does not adhere to operator preference. se-r does not address viewpoint invariance which is a significant problem in vision-based off-road navigation and hence performs poorly in Envs. 1 and 2. We perform additional experiments in an outdoor environment (Env. 6) to study adherence to operator preferences, detailed in Supplementary Sec. 9.1.
@@ 6 Conclusion
In this paper, we introduce Self-supervised TErrain Representation LearnING (sterling), a novel framework for learning terrain representations from easy-to-collect, unconstrained (e.g., non-expert), and unlabeled robot experience. sterling utilizes non-contrastive representation learning through viewpoint invariance and multi-modal correlation self-supervision objectives to learn relevant terrain representations for visual navigation. We show how features learned through sterling can be utilized to learn operator preferences over terrains and integrated within a planner for preference-aligned navigation. We evaluate sterling against state-of-the-art alternatives on the task of preference-aligned visual navigation on a Spot robot and find that sterling outperforms other methods and performs on par with a fully-supervised baseline. We additionally perform a qualitative large-scale experiment by successfully hiking a 3-mile-long trail using sterling, demonstrating its robustness to off-road conditions in the real world.